A U.S. census bureau pollster noted that in 358 random households surveyed, 233 occupants owned their own home. For a 95% confidence interval for the proportion of home owners, what is the margin of error?

Question 4 options:

1)

0.0413

2)

0.0494

3)

0.0012

4)

0.0026

5)

0.0252
Answer #1

Solution :

Given that,

n = 358

x =233

Point estimate = sample proportion = \hat p = x / n = 233/358=0.651

1 - \hat p = 1 - 0.651=0.349

At 95% confidence level

\alpha = 1 - 95%  

\alpha = 1 - 0.95 =0.05

\alpha/2 = 0.025

Z\alpha/2 = Z0.025 = 1.96

Margin of error = E = Z\alpha / 2 * \sqrt ((\hat p * (1 - \hat p )) / n)

= 1.96 (\sqrt((0.651*0.349) / 358)

E = 0.0494

correct option is

2)

0.0494